# On the Entanglement Entropy Distribution of a Hybrid Quantum Circuit

### 31 Mar 2026 v1

@@ Abstract
We investigate the distribution of entanglement entropy in hybrid quantum circuits consisting of random unitary gates and local measurements applied at a finite rate. We demonstrate that higher moments of the entanglement entropy distribution, such as the ratio of the variance to the mean and the skewness, capture nontrivial features of the measurement-induced dynamics that are invisible to the mean entropy alone. We demonstrate that these quantities exhibit distinct and robust behaviors across the volume-law and area-law phases, and can serve as effective diagnostics of measurement-induced entanglement transitions. We propose a phenomenological model describing the effect of measurements in the area-law regime, which, when combined with the directed polymer in a random environment description of the volume-law phase, matches numerical simulations well across the entire phase diagram.
@@ I Introduction
Random unitary circuits (RUCs) provide a minimal model that extends our understanding of the chaotic evolution of quantum many-body systems [12, 40, 18, 52]. Accompanied by chaotic scrambling of information, random unitary evolution results in an extensive scaling of the entanglement entropy of the system [37, 38, 61]. Contrary to unitary evolution, measurement on a quantum system is non-unitary and irreversible. The peculiar properties of measurement give rise to rich phenomena in various quantum systems [57]. In quantum information science, measurement-based quantum computation utilizes such properties [43, 8]. In many-body physics, the quantum Zeno effect demonstrates that frequent measurements result in an area-law scaling of the entanglement entropy of the system [35, 29].
@@
MIPT has been extensively studied in various fields of physics. From the original quantum circuits which were explored [29, 48, 30, 21, 36, 12, 49, 19, 36], it has been extended to a wide range of many-body systems such as free fermionic systems [42, 41, 34], Ising-like models [28, 51], and bosonic systems [59]. In the context of quantum information science, MIPT has been extended to information exchanging processes [23]. In understanding the entanglement dynamics of the hybrid circuit, the non-unitarity of the measurements poses a significant challenge to the analysis of the transition. While mappings of hybrid quantum circuits to statistical-mechanics models have successfully yielded scaling properties [4, 21] and provided an exact mapping in the volume-law phase [31], a complete understanding of the entanglement behavior across the entire range of measurement probabilities remains open, especially in the area-law phase.
@@
The MIPT is usually characterized by the scaling of the mean entanglement entropy, with only a few works examining higher moments such as the variance [49, 24]. However, hybrid quantum circuits are intrinsically non-self-averaging due to the stochastic nature of measurements. As a result, the entanglement entropy exhibits strong trajectory-to-trajectory fluctuations that persist even in the thermodynamic limit [3, 44]. Moreover, in a directed polymer in a random environment (DPRE) [31], which provides an effective description of hybrid circuits in the volume-law phase, the entanglement entropy plays the role of a fluctuating free energy, whose higher moments encode universal features of the underlying stochastic dynamics [58]. In such a situation, the average entanglement entropy captures only the leading extensive contribution, while finer statistical structures could be hidden in higher-order moments.
In this work, we explore higher-order moments of the entanglement entropy distributions in hybrid quantum circuits beyond their mean values. We first introduce the index of dispersion (IoD), defined as the ratio of the variance to the mean, and demonstrate that it exhibits qualitatively distinct behaviors in the volume-law and area-law phases. We further observe an abrupt change in the IoD near the critical point, enabling a precise estimate of $p_{c}$, which cannot be obtained from the variance alone. We further analyze a higher-order moment of the entanglement entropy distribution by examining its skewness, which quantifies the distribution’s asymmetry using moments up to third order. Remarkably, independent of the system size, the skewness remains constant in the volume law regime, while it follows power-law scaling in the area law regime. Similar to the IoD, the skewness increases sharply near the critical probability, providing a sensitive and robust diagnostic of MIPT in the hybrid circuit.
@@ II.2 Average entanglement and MIPT
In the regime where the rate of measurement is sufficiently low, i.e., $p<p_{c}$, quantum correlations in the system prevail, so that the average entanglement entropy of the system follows the volume law scaling, $\langle S\rangle\propto L$. On the other hand, frequent measurements on the system, i.e., $p>p_{c}$, prevents correlations among the qubits from forming, resulting in the area law scaling of the entanglement entropy, $\langle S\rangle\propto L^{0}$.
@@ III Higher-order moments of the entanglement distribution
Figure 5: Variance curves of both Haar random and Clifford random hybrid quantum circuits. The critical measurement rate for the Haar (Clifford) random circuit is indicated by a solid (dashed) vertical line. $p_{c}=0.26$ for Haar random circuit and $p_{c}=0.16$ for Clifford random circuit.
@@ III.1 Variance
$\mathrm{Var}[S]=\langle(S-\langle S\rangle)^{2}\rangle$ (6)
@@
In particular, the peak of the variance ($p_{\max}$) for large system sizes deviates from $p_{c}$ and lies within the volume law phase. Moreover, the variance exhibits a sharp peak deep within the volume-law regime, rendering the precise location of the maximum ambiguous. These observations support the conclusion that variance alone may not provide sufficient evidence for identifying the critical point of the hybrid quantum circuit.
@@
Figure 6: Index of dispersion (IoD) for varying rate of measurement. The two distinct phases are clearly observed in the IoD curves. The inset shows that the discontinuity in the curves occurs near the critical point, $p_{c}=0.16$ (dashed vertical gray line).
@@ III.2 Index of dispersion
The observation that variance alone may be insufficient to fully characterize MIPT motivates us to develop a new measure that more sensitively reflects changes in the entanglement distribution with respect to the measurement rate. As an illustrative example, when the entanglement entropy is bounded by a condition $S\geq 0$, the variance of the entanglement distribution in the area law regime is bounded by $\mathrm{Var}[S]\sim L^{0}$. This implies that considering the variance rescaled by the mean value would be more appropriate in accounting for the change in entanglement distribution.
@@
As shown in the Fig. 6, we find that the IoD exhibits qualitatively distinct behavior across the two entanglement phases. In the volume-law regime, the IoD curves vary with different system sizes with the same rate of measurement. Whereas in the area-law regime, the curves of different system sizes converge to a single curve. As the measurement rate increases, the IoD approaches unity, indicating that the entanglement entropy distribution becomes approximately Poissonian. The inset of the Fig. 6 shows the discontinuity between the volume-law regime and area-law regime which occurs at the critical point, suggesting the possible use of the IoD to detect the phase transition.
@@ III.3 Skewness
While the IoD can efficiently indicate the critical behavior of MIPT using only the mean and variance of the entanglement distribution, these up to second-moment quantities do not capture the detailed shape of the distribution, for example, symmetry around the mean value. To explore a more detailed distribution of the entanglement entropy, we utilize the skewness, the lowest-order moment sensitive to the asymmetry of a given distribution. It is used in various areas of physics, including fluid dynamics [60], plasma physics [15], and statistical physics [54, 47]. Especially in the context of statistical physics, higher moments of fluctuating quantities can be used to distinguish the universality class of the system [47].
@@
Figure 7 shows skewness curves for different system sizes. In the volume law regime, despite the wide range of system sizes, the skewness remains at a constant value of $\gamma(S_{\mathrm{Vol}})=-0.224$ regardless of the system sizes. This constant value results from a previously known mapping between a hybrid quantum circuit in the volume-law regime and a statistical-mechanics model [31]. We defer the details on this matter to the later part of this work. The invariance of the skewness in the volume law regime distinguishes it from other moments, such as the mean, variance, or IoD. On the other side of the phase transition, the skewness curve gradually increases in the area law phase. This behavior becomes evident in a log-scaled plot as shown in Figure 8, revealing a power-law scaling with $p$, $\gamma(S)=7.10p^{1.79}$.
@@
Between the two scaling phases, a sudden increase in skewness is observed, indicating significant deformation of the distribution shape at the transition between the phases. To address a point where the change in skewness of distributions is the most significant, we employ a smooth auxiliary function to locate the local maximum. The choice of auxiliary function should be chosen such that it describes the transition between the distinct behaviors of skewness in the two scaling phases near the criticality:
@@
Taking into account the power law scaling in the area law regime, a smooth function was fitted to the skewness data for $L=512$ and $1024$ to estimate the point of the maximal change in skewness. We take the following form of the auxiliary function,
@@
where $b_{0,1,2}$ are constants obtained by minimizing mean-squared distances between the data and the curve for each system sizes. The first term captures the asymptotic power-law scaling in the area-law phase, whereas the latter term is introduced to model a crossover near the critical point. The hyperbolic tangent function provides a smooth interpolation between two asymptotic regimes and is therefore well-suited to describe a localized but continuous change in the slope of the skewness curve. In this manner, the function introduced in Eq. (10) serves as a phenomenological tool to robustly identify the point of maximal deformation in the entanglement entropy distribution. In Fig. 9, we present fits of the auxiliary function to our simulation results, which agree well with the data.
@@ IV.1 Volume law phase
$\mathrm{IoD}=\frac{s_{1}^{2}L_{A}^{1/3}\mathrm{Var}[\xi]}{s_{0}L_{A}^{2/3}-s_{1}\langle\xi\rangle}$ (12)
@@
$=-\gamma(\xi)$ (14)
where $\gamma(\xi)=-0.224$ [9], in agreement with our numerical results in the volume-law regime. Therefore, it provides independent numerical confirmation of the DPRE description established in Ref. [31] for the monitored quantum circuits with periodic boundary conditions.
IV.2 Area law phase
To describe the entanglement entropy in the area law, we introduce the following stochastic model by assuming a simple structure for the entanglement entropy and its change after saturation. We consider the change in entanglement entropy over a single time step:
@@ IV.2 Area law phase
where $\sigma_{U(M)}$ is a random variable representing the effect of unitary (measurement) operations, which generally increase (decrease) the entanglement entropy. For simplicity, we assume that $\sigma_{U}$ follows a Gaussian distribution with mean $\mu$ and variance of $\nu$, both of which are determined numerically from simulations involving only unitary gates.
@@
$r=\frac{\binom{L}{N}-2\binom{L-2}{N}+\binom{L-4}{N}}{\binom{L}{N}}$. (18)
and the variance is given by $q(1-q)S+(r-q^{2})S(S-1)$ where $r-q^{2}$ is the covariance of the event of two distinct pairs are destroyed. $r-q^{2}$ can be calculated algebraically and in the large $L$ limit, $r-q^{2}=-\frac{4p(1-p)^{3}}{L}+O(L^{-2})$. However, we found that the calculated value does not accurately describe the numerical data. Hence, we replaced $r-q^{2}$ with $c$, a quadratic function of $q$, whose coefficients are optimized using our numerical data for each $p$.
@@
$b(S)\equiv\mathrm{Var}(\Delta S|S)$ $=\nu+q(1-q)S+cS(S-1).$ (20)
@@
$\partial_{t}P(S,t)=-\partial_{S}[aP(S,t)]+\frac{1}{2}\partial^{2}_{S}[bP(S,t)].$ (21)
The steady state solution, $P(S)$, satisfies $\partial_{t}P(S)=0$, resulting in the ordinary differential equation
@@
Solving the equation gives the stationary distribution $P(S)$,
@@
$\propto\frac{1}{b}\exp(\int\frac{2(\mu-qS)}{\nu+S\alpha+S(S-1)c}dS),$ (24)
where $b^{\prime}=\frac{db}{dS}$ and $\alpha=q(1-q)$. Since $c$ is small, the integral in the exponent can be expanded perturbatively up to second order in $c$. This yields the entanglement distribution $P(S)\propto\frac{1}{b}P_{0}P_{1}P_{2}$ where $P_{i}$ denotes the contribution at order $c^{i}$:
$P_{0}$ $=x^{\gamma}e^{-\frac{2s}{1-q}}$ (25)
$P_{1}$ $=x^{d_{1}c}\exp[-c\{d_{0}x^{-1}+d_{2}x+\frac{d_{3}}{2}x^{2}\}]$ (26)
$P_{2}$ $=x^{e_{1}c^{2}}\exp[-c^{2}\{e_{o}x^{-1}+e_{2}x+\frac{e_{3}}{2}x^{2}+\frac{e_{4}}{3}x^{3}\}]$ (27)
where $x=(\nu+\alpha S)$ and $\gamma=\frac{2}{\alpha}(\frac{\nu}{1-q}+\mu)$. The coefficients $d_{i}$ and $e_{i}$ are constants determined by the perturbative expansion. Although the resulting expression for $P(S)$ is algebraically involved, once the conditional moments are specified the remaining steps follow straightforwardly.
We emphasize that this derivation of the Fokker-Planck equation is a coarse grained stochastic description, not a fundamental equation of motion. Therefore, the equation neglects the microscopic details of the underlying circuit structure and dynamics. However, this approach allows for a low dimensional description to capture the entanglement entropy distribution in the area law phase.

### 13 Apr 2026 v2

@@ Abstract
We investigate the distribution of entanglement entropy in hybrid quantum circuits consisting of random unitary gates and local measurements applied at a finite rate. We demonstrate that higher moments of the entanglement entropy distribution, such as the ratio of the variance to the mean and the skewness, capture nontrivial features of the measurement-induced dynamics that are invisible to the mean entropy alone. We demonstrate that these quantities exhibit distinct and robust behaviors across the volume-law and area-law phases, and can serve as effective diagnostics of measurement-induced entanglement transitions. We propose a phenomenological model describing the effect of measurements in the area-law regime, which, when combined with the directed polymer in a random environment description of the volume-law phase, matches numerical simulations well across both phases.
@@ I Introduction
Random unitary circuits (RUCs) provide a simple paradigmatic model that extends our understanding of the chaotic evolution of quantum many-body systems [13, 41, 19, 54]. Accompanied by chaotic scrambling of information, random unitary evolution results in an extensive scaling of the entanglement entropy of the system [38, 39, 63]. Contrary to unitary evolution, measurement on a quantum system is non-unitary and irreversible. The peculiar properties of measurement give rise to rich phenomena in various quantum systems [59]. In quantum information science, measurement-based quantum computation utilizes such properties [44, 9]. In many-body physics, the quantum Zeno effect demonstrates that frequent measurements result in an area-law scaling of the entanglement entropy of the system [36, 30].
@@
MIPT has been extensively studied in various fields of physics. Since the first proposal [30] consisting of random unitary circuits with local measurements [30, 50, 31, 22, 37, 13, 51, 20, 37], such a formalism has been extended to a wide range of many-body systems such as free fermionic systems [43, 42, 35], Ising-like models [29, 53], and bosonic systems [61]. In the context of quantum information science, MIPT has been extended to information exchange processes [24]. In understanding the entanglement dynamics of the hybrid circuit, the non-unitarity of the measurements poses a significant challenge to the analysis of the transition. While mappings of hybrid quantum circuits to statistical-mechanics models have successfully yielded scaling properties [4, 22] and provided an exact mapping in the volume-law phase [32], a complete understanding of the entanglement behavior across the entire range of measurement probabilities remains open, especially in the area-law phase.
@@
The MIPT is usually characterized by the scaling of the mean entanglement entropy, with only a few works examining higher moments such as variance [51, 25]. However, hybrid quantum circuits are intrinsically non-self-averaging due to the stochastic nature of measurements. As a result, the entanglement entropy exhibits strong trajectory-to-trajectory fluctuations that persist even in the thermodynamic limit [3, 46]. Moreover, in a directed polymer in a random environment (DPRE) [32], which provides an effective description of hybrid circuits in the volume-law phase, the entanglement entropy plays the role of a fluctuating free energy, whose higher moments encode universal features of the underlying stochastic dynamics [60]. In such a situation, the average entanglement entropy captures only the leading extensive contribution, while finer statistical structures could be hidden in higher-order moments.
In this work, we explore higher-order moments of the entanglement entropy distributions in hybrid quantum circuits beyond their mean values. We first introduce the index of dispersion (IoD), defined as the ratio of the variance to the mean, and demonstrate that it exhibits qualitatively distinct behaviors in the volume-law and area-law phases. We further observe an abrupt change in the IoD near the critical point, enabling a precise estimate of $p_{c}$, which cannot be obtained from the variance alone. We further analyze a higher-order moment of the entanglement entropy distribution by examining its skewness, which quantifies the distribution’s asymmetry using moments up to third order. Remarkably, independent of the system size, the skewness remains constant in the volume law regime, while it follows power-law scaling in the area-law regime. Similar to the IoD, the skewness increases sharply near the critical probability, providing a sensitive and robust diagnostic of MIPT in the hybrid circuit.
@@ II.2 Average entanglement and MIPT
In the regime where the rate of measurement is sufficiently low, i.e., $p<p_{c}$, quantum correlations in the system prevail, so that the average entanglement entropy of the system follows the volume law scaling, $\langle S\rangle\propto L$. On the other hand, frequent measurements on the system, i.e., $p>p_{c}$, prevents correlations among the qubits from forming, resulting in the area-law scaling of the entanglement entropy, $\langle S\rangle\propto L^{0}$.
@@ III Higher-order moments of the entanglement distribution
Figure 5: Variance of Haar random and Clifford random hybrid quantum circuits. The triangle (circle) points indicate the peak of the variance for the Haar (Clifford) random circuit. The critical measurement rate for the Haar (Clifford) random circuit is indicated by a solid (dashed) vertical line. $p_{c}=0.26$ for Haar random circuit and $p_{c}=0.16$ for Clifford random circuit.
@@ III.1 Variance
$\mathrm{Var}[S]=\langle(S-\langle S\rangle)^{2}\rangle,$ (6)
@@
In particular, the peak of the variance for large system sizes deviates from $p_{c}$ and lies within the volume law phase. Moreover, the variance exhibits a sharp peak deep within the volume-law regime, rendering the precise location of the maximum ambiguous. These observations support the conclusion that variance alone may not provide sufficient evidence for identifying the critical point of the hybrid quantum circuit.
@@ III.2 Index of dispersion
The observation that variance alone may be insufficient to fully characterize MIPT motivates us to develop a new measure that more sensitively reflects changes in the entanglement distribution with respect to the measurement rate. As an illustrative example, when the entanglement entropy is bounded by a condition $S\geq 0$, the variance of the entanglement distribution in the area-law regime is bounded by $\mathrm{Var}[S]\sim L^{0}$. This implies that considering the variance rescaled by the mean value would be more appropriate in accounting for the change in entanglement distribution.
@@
Figure 6: Index of dispersion (IoD) for varying rate of measurement. The two distinct phases are clearly observed in the IoD curves. The inset shows that the discontinuity in the curves occurs near the critical point, $p_{c}=0.16$ (dashed vertical line).
As shown in the Fig. 6, we find that the IoD exhibits qualitatively distinct behavior across the two entanglement phases. In the volume-law regime, the IoD curves vary with different system sizes at the same measurement rate. Whereas in the area-law regime, the curves with different system sizes converge to a single curve. As the measurement rate increases, the IoD approaches unity, indicating that the entanglement entropy distribution becomes approximately Poissonian. The inset of Fig. 6 shows the discontinuity between the volume-law regime and area-law regime, which occurs at the critical point, suggesting the possible use of the IoD to detect the phase transition.
@@ III.3 Skewness
While the IoD can efficiently indicate the critical behavior of MIPT using only the mean and variance of the entanglement distribution, these up to second-moment quantities do not capture the detailed shape of the distribution, for example, symmetry around the mean value. To explore a more detailed distribution of the entanglement entropy, we utilize the skewness, the lowest-order moment sensitive to the asymmetry of a given distribution. It is used in various areas of physics, including fluid dynamics [62], plasma physics [16], and statistical physics [56, 49]. Especially in the context of statistical physics, higher-order moments of fluctuating quantities can be used to distinguish the universality class of the system [49].
@@
Figure 7 shows skewness curves for different system sizes. In the volume law regime, despite the wide range of system sizes, the skewness remains at a constant value of $\gamma(S_{\mathrm{Vol}})=-0.224$ regardless of the system sizes. This constant value results from a previously known mapping between a hybrid quantum circuit in the volume-law regime and a statistical-mechanics model [32]. We defer the details on this matter to the latter part of this work. The invariance of skewness in the volume-law regime distinguishes it from other moments, such as the mean, variance, or IoD. On the other side of the phase transition, the skewness curve gradually increases in the area-law phase. This behavior becomes evident in a log-scaled plot as shown in Fig. 8, revealing a power-law scaling with $p$, $\gamma(S)=7.10p^{1.79}$.
@@
Between the two scaling phases, a sudden increase in skewness is observed, indicating significant deformation of the entanglement entropy distribution at the transition point. To address the point at which the change in skewness of distributions becomes the most significant, we employ a smooth auxiliary function to locate the local maximum. The choice of auxiliary function should be taken such that it describes the transition between the distinct behaviors of skewness in the two scaling phases near the criticality as follows:
@@
Taking into account the power law scaling in the area-law regime, a smooth function was fitted to the skewness data for $L=512$ and $1024$ to estimate the point of the maximal change in skewness. We take the following form of the auxiliary function,
@@
where $b_{0,1,2}$ are constants obtained by minimizing mean-squared distances between the data and the curve for each system size. The first term captures the asymptotic power-law scaling in the area-law phase, whereas the latter term is introduced to model a crossover near the critical point. The hyperbolic tangent function provides a smooth interpolation between two asymptotic regimes and is therefore well-suited to describe a localized but continuous change in the slope of the skewness curve. In this manner, the function introduced in Eq. (10) serves as a phenomenological tool to robustly identify the point of maximal deformation in the entanglement entropy distribution. In Fig. 9, we present fits of the auxiliary function to our simulation results, which agree well with the data.
@@ IV.1 Volume law phase
$\mathrm{IoD}=\frac{s_{1}^{2}L_{A}^{1/3}\mathrm{Var}[\xi]}{s_{0}L_{A}^{2/3}-s_{1}\langle\xi\rangle},$ (12)
@@
$=-\gamma(\xi),$ (14)
where $\gamma(\xi)=0.224$ [10], in agreement with our numerical results in the volume-law regime. Therefore, it provides independent numerical confirmation of the DPRE description established in Ref. [32] for the monitored quantum circuits with periodic boundary conditions.
IV.2 Area-law phase
To describe the entanglement entropy in the area-law, we introduce the following stochastic model by assuming a simple structure for the entanglement entropy and its change after saturation. We consider the change in entanglement entropy over a single time step:
@@ IV.2 Area-law phase
where $\sigma_{U(M)}$ is a random variable representing the effect of unitary (measurement) operations, which generally increase (decrease) the entanglement entropy. For simplicity, we assume that $\sigma_{U}$ follows a Gaussian distribution with mean $\mu=0.58$ and variance of $\nu=0.80$, both of which are determined numerically from simulations.
@@
$r=\frac{\binom{L}{N}-2\binom{L-2}{N}+\binom{L-4}{N}}{\binom{L}{N}}$, (18)
and the variance is given by $q(1-q)S+(r-q^{2})S(S-1)$, where $r-q^{2}$ is the covariance of the event that two distinct pairs are destroyed. $r-q^{2}$ can be calculated algebraically and in the large $L$ limit, $r-q^{2}\xrightarrow{L\gg 1}-\frac{4p(1-p)^{3}}{L}+O(L^{-2})$. However, this form does not accurately describe the numerical data, as the correlation is strongly dependent on $q$, and we replaced $r-q^{2}$ with $c(q)$, a quadratic function of $q$ whose coefficients are optimized from our numerical data for each $p$.
@@
$b(S)\equiv\mathrm{Var}(\Delta S|S)$ $=\nu+q(1-q)S+c(q)S(S-1).$ (20)
@@
$\partial_{t}P(S,t)$ $=-\partial_{S}[aP(S,t)]+\frac{1}{2}\partial^{2}_{S}[bP(S,t)]$ (21)
$=-\partial_{S}J(S,t)$ (22)
where the second line is the continuity equation with probability current
$J(S,t)=a(S)P(S,t)-\frac{1}{2}\partial_{S}[b(S)P(S,t)].$ (23)
In the stationary state, $\partial_{t}P(S)=0$, so that $\partial_{S}J(S)=0$ and hence the current is independent of $S$. Imposing the zero-flux boundary condition at the boundaries of the allowed domain of $S$, ($0\leq S\leq L/2$), we obtain $J(S)=0$, which yields
@@
Solving the equation gives the stationary distribution $P(S)$ [45],
@@
$\propto\frac{1}{b}\exp(\int\frac{2(\mu-qS)}{\nu+\alpha S+c(q)S(S-1)}dS),$ (26)
where $b^{\prime}=\frac{db}{dS}$ and $\alpha=q(1-q)$. Since the covariance $c(q)$ is small, in general, we treat it perturbatively and perform a Maclaurin expansion of the integrand up to second order in $c(q)$ (see Appendix A for more details).
We emphasize that this derivation of the Fokker-Planck equation is a coarse-grained stochastic description, not a fundamental equation of motion. Therefore, the equation neglects the microscopic details of the underlying circuit structure and dynamics. However, this approach allows a low-dimensional description to capture the entanglement entropy distribution in the area-law phase.